A company estimates that the demand for its product fluctuates with the price it charges. The demand function is given as:
q = 100,000 - 200p

Where .q. equals the number of units demanded and .p. equals the price in rupees. The total cost of producing q units of the product is estimated by the function:

C = 150,000 + 100q + 0.003q2
Required:

a.Determine how many units of q should be produced in order to maximize annual profit?
b.What price should be charged?
c.What is the annual profit expected to equal?
1
Expert's answer
2020-07-30T15:29:35-0400
"Profit=Revenue-Cost"

"q = 100000 - 200p=>p(q)=500-0.005q"

"R(q)=p(q)\cdot q=500q-0.005q^2"

"P(q)=R(q)-C(q)=""=500q-0.005q^2-(150000 + 100q + 0.003q^2)=""=400q-0.008q^2-150000"

a.

"P'(q)=400-0.016q"

Find the critical number(s)

"P'(q)=0=>400-0.016q=0=>q=25000"

"P''(q)=-0.016<0"

The function "P(q)" has a local maximum at "q=25000."

Since the function "P(q)" has the only extremum, then the function "P(q)" has the absolute maximum at "q=25000."

A company has to produce "q=25000" units in order to maximize annual profit.


b.

"p(25000)=500-0.005(25000)=375"

The price is Rs375.

c.


"P(25000)=400(25000)-0.008(25000)^2-150000="

"=4850000"

The expected annual profit is Rs4,850,000.